OpenStudy (anonymous):

What is an equation for the line passes your point (8,9) and (10,3) ? A) (y -9)=-3(x-8) B) (y +9)=-3(x+8) C) (y -9)= 3(x-8) D) (y +9)=3(x+8)

OpenStudy (anonymous):

well since two points on the line are given you can first find the slope of the line with the slope formula m=y2-y1/x2-x1. you can then utilize the first point, 8,9 to plug into the y-y1=(m{slope})(x-x1) and that will give you your answer. :)

OpenStudy (anonymous):

I did the first part n got -3 but then I'm confused

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

Well, assuming -3 is the correct slope, you would just plug in y-9=-3(x-8) which then becomes y-9=-3x+24 (after distribution) -9 -9 then your going to minus 9 from both sides y=-3x+15 ok

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

no problem.